### v2

# Quality Time: Carbon-Aware Quality Adaptation for Energy-Intensive Services

Abstract.
The energy demand of modern cloud services, particularly those related to generative AI, is increasing at an unprecedented pace. While hyperscalers collectively fail to meet their self-imposed emission reduction targets, they face increasing pressure from environmental sustainability reporting across many jurisdictions. To date, carbon-aware computing strategies have primarily focused on batch process scheduling or geo-distributed load balancing. However, such approaches are not applicable to services that require constant availability at specific locations due to latency, privacy, data, or infrastructure constraints.
In this paper, we explore how the carbon footprint of energy-intensive services can be reduced by adjusting the fraction of requests served by different service quality tiers. We show that adapting this quality of responses with respect to grid carbon intensity can lead to additional carbon savings beyond resource and energy efficiency. Building on this, we introduce a forecast-based multi-horizon optimization that reaches close-to-optimal carbon savings and is able to automatically adapt service quality for best-effort users to stay within an annual carbon budget. Our approach can reduce the emissions of large-scale LLM services, which we estimate at multiple 10,000 tons of CO2 annually, by up to 10 %.
Sustainable computing, quality of service, LLM inference, green AI
submissionid: ¡PAPER ID¿ccs: Social and professional topics Sustainabilityccs: Networks Cloud computingPreprint v2. Under review.
1. Introduction
Operating modern cloud services can require substantial computing resources and energy, which directly contributes to the rapidly growing carbon footprint of computing systems (S&P Globl Market Intelligence, 2023). Alongside traditional resource-intensive services like real-time data analytics, video processing, recommendation systems, and streaming, the growing use and high energy demand of generative AI is currently driving the expansion of data centers at a high pace (Bashir et al., 2024; Wu et al., 2024). As a result, major cloud providers, who collectively pledged to reduce their carbon footprint to zero by 2030, have been increasing their emissions significantly in recent years (Google, 2024; Microsoft, 2024; Meta Platforms, 2023) and are expected to row back on sustainability commitments (Lawrence, 2024).
Figure 1. The carbon intensity of electricity can vary significantly over time. Exploiting these temporal variations can reduce the carbon footprint of energy-intensive services.
Besides their social responsibility to minimize environmental impact, large companies are increasingly under pressure to report their greenhouse gas (GHG) emissions as part of environmental sustainability reporting (Project, 2024; initiative, SBTi). While reporting on Scope 2 emissions (GHG emissions caused by electricity consumption) is becoming mandatory in many parts of the world (Deloitte, 2024; European Parliament and Council of the European Union, 2022; UK Department for Business, Energy & Industrial Strategy, 2019; Japan Ministry of Enconomy, Trade and Industry (2023), METI), upcoming policy changes will mandate that large companies also disclose Scope 3 emissions (European Parliament and Council of the European Union, 2022; Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023)—including emissions from the use of cloud services. This will affect not only service providers, but also users of energy-intensive services.
The concept of carbon-aware computing has emerged as a response to such challenges: Carbon-aware strategies aim to reduce emissions beyond energy efficiency by additionally taking into account how carbon-intensive the consumption of electricity at a current time and location is. Especially in the context of scheduling flexible batch workloads, the potential of temporal and spatial load shifting is well understood (Wiesner et al., 2021; Radovanovic et al., 2022; Lechowicz et al., 2023; Zheng et al., 2020; Wiesner et al., 2024; Hanafy et al., 2024b; Sukprasert et al., 2024b; Hanafy et al., 2024a). However, for continuously running services, where individual requests have little to no delay-tolerance, existing approaches exclusively focus on geographical load balancing (Gsteiger et al., 2024; Murillo et al., 2024; Souza et al., 2023; Zhou et al., 2016, 2013). However, many services cannot simply be distributed across regions due to data locality requirements, regulatory constraints, privacy and security concerns, or infrastructure limitations. For instance, generative AI inference typically relies on high-performance GPUs or specialized accelerators, which are often only available at specific locations. Carbon-aware approaches for services that are constrained to a single region remain largely unexplored.
Figure 2. We investigate how adjusting the QoR with respect to carbon intensity can reduce emissions beyond savings related to reduced energy demand.
In this work, we focus on the temporal variability of carbon intensity (see Figure 1) and explain how adapting service quality with respect to this metric does not only impact energy demand, but can yield additional carbon savings. We define the quality of responses (QoR) as a type of quality of service (QoS) metric that describes the proportion of requests served by different service quality tiers. For instance, machine learning models can be deployed in smaller, quantized, or pruned versions that use fewer resources and consume less energy, but compromise on response quality (Liang et al., 2021). Our approach minimizes the operational carbon emissions of energy-intensive services by optimizing resource provisioning and load balancing across service quality tiers, while ensuring compliance with service level objectives (SLOs) for different user groups.
Figure 2 shows the accumulated energy usage and carbon emissions for an example scenario with two quality tiers over one week. Serving all requests at the high quality tier (QoR = 1) uses roughly twice the energy compared to serving them at the low quality tier (QoR = 0). Balancing requests evenly between the two tiers results in intermediate energy usage. We explore how strategically timing quality reductions based on carbon intensity can further reduce emissions by performing a comprehensive analysis across different regions, request patterns, and QoR constraints. Based on this, we propose an approach that achieves close-to-optimal savings despite real-life challenges such as forecast inaccuracies and the high computational complexity of the underlying optimization.
Contributions.  Towards more carbon-efficient cloud services, we make the following contributions:
  1. (1)
    We formalize the problem of minimizing carbon emissions under QoR constraints through optimized resource provisioning and load balancing between service quality tiers.
  2. (2)
    We model emissions from the perspective of data center operators and cloud tenants (Scope 2 and 3 reporting) and prove that both models lead to equivalent decisions.
  3. (3)
    We propose a forecast-based multi-horizon optimization for minimizing service emissions, which can automatically adjust the QoR for best-effort users to stay within a predefined annual carbon budget.
  4. (4)
    We evaluate the upper-bound potential and practicability of carbon-aware QoR adaptation by simulating a large-scale LLM inference service on eight real and synthetic request traces in ten different regions.
All experiments in this paper are fully reproducible. All code and data will be made publicly available upon acceptance.
2. Background and Motivation
Despite previous commitments to achieve fully carbon-free or even “carbon-negative” operations by 2030, Microsoft recently announced a 29.1 % rise in emissions since 2020 (Microsoft, 2024), Google reported a 48 % increase in GHG emissions within five years (Google, 2024), while Meta increased their emissions by 46 % in 2022 alone (Meta Platforms, 2023)—predominantly because of the training and deployment of generative AI. The rapid expansion of data centers, combined with the slow adoption of clean energy and efficiency measures, suggests that reaching net-zero emissions in the near future is highly unlikely (Lawrence, 2024).
All big hyperscalers have announced or introduced carbon reporting tools for their customers (Google Cloud, 2023; Microsoft, 2023; Amazon Web Services, 2023; Dixit and Tse, 2024). However, they are expected to significantly underreport emissions, due to opaque and inconsistent reporting methodologies (O’Brien, 2024). We anticipate an increasing focus on regulatory frameworks that will require cloud providers to disclose their carbon emissions more comprehensively.
Energy-intensive cloud services.  The rapid adoption of large language models (LLMs) requires immense processing power, not only for training but also for inference at scale (Faiz et al., 2024; Luccioni et al., 2024). As generative AI evolves towards multi-modality, emerging formats like video are expected to further increase future electricity demand (Li et al., 2024a).
Beyond generative AI, a range of current and emerging services are contributing to increasing energy demands. Video streaming, predominantly in high-resolution formats, now accounts for 82 % of all internet traffic (Systems, 2024). Platforms such as cloud gaming, virtual reality, and the metaverse require real-time graphics rendering, which consumes significant power (Stoll et al., 2022; Mills et al., 2019; Kattakinda et al., 2024). Additionally, scientific simulations such as climate modeling (Acosta et al., 2024), drug discovery (Wynendaele et al., 2021), astrophysics (Taffoni et al., 2019), fluid dynamics (Yang et al., 2024), and bioinformatics (Grealey et al., 2022) are growing in complexity and increasingly integrate AI services.
Growing pressure from GHG reporting.  The reporting of Scope 2 emissions, i.e. indirect GHG emissions caused by consuming electricity, is becoming increasingly standardized and mandatory across major jurisdictions (Deloitte, 2024; European Parliament and Council of the European Union, 2022; UK Department for Business, Energy & Industrial Strategy, 2019; Japan Ministry of Enconomy, Trade and Industry (2023), METI; Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023), with significant implications for data center operators, cloud tenants, and—ultimately—users of energy-intensive services.
In the US, the Securities and Exchange Commission (SEC) is advancing regulations that require public companies to fully disclose their Scope 2 emissions (Deloitte, 2024). While, the disclosure of Scope 3 emissions—i.e. all indirect emissions that occur across the value chain—stays voluntarily, for now, California soon requires large companies to also report Scope 3 emissions (Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023). Meanwhile, the EU’s Corporate Sustainability Reporting Directive (CSRD) mandates Scope 2 and 3 disclosures for large companies (European Parliament and Council of the European Union, 2022) beginning in 2024, aligning with the EU’s broader climate goals under the European Green Deal (European Commission, 2019). From 2026 on, this will also include non-EU companies with significant operations in the EU. As the significance of Scope 3 emissions has been emphasized repeatedly (Hettler and Graf‐Vlachy, 2024; Hertwich and Wood, 2018; Radonjič and Tompa, 2018), also jurisdictions outside the EU and California are likely to require Scope 3 emissions reporting in the future.
Implications for energy-intensive services.  When reporting emissions from the perspective of cloud service providers, we must distinguish between two cases:
  1. (1)
    They operate data centers themselves, in which case the operational carbon footprint is mainly determined by electricity usage, i.e. Scope 2 emissions.
  2. (2)
    They rent infrastructure from cloud providers, in which case the data center’s emissions will be passed down the value chain als Scope 3 emissions, which is an active topic of research (Amaral et al., 2023; Choochotkaew et al., 2023; Hè et al., 2023; Schmidt et al., 2023; Westerhof et al., 2023; Han et al., 2024).
Today’s, landscape is mixed: While, for example, Google relies on its proprietary infrastructure to train and serve AI models like Gemini (Pichai and Hassabis, 2024), Microsoft Azure provides the computational resources for operating OpenAI’s ChatGPT (OpenAI, 2023). We will address the two cases in our problem formalization, however, both are affected by stricter GHG reporting. Following, the increasing emissions of computing and the growing pressure for comprehensive accounting will soon have a direct impact on service cost. Carbon pricing mechanisms covered 23 % of global emissions in 2023, where revenues reached a record \$104 billion, and there is a clear trend towards stricter rules and higher prices (World Bank, 2024).
Assumptions on future regulatory policies.  As the effectiveness of annual and globally valid energy certificates is highly questionable (Bjørn et al., 2022; Holzapfel et al., 2023; Bashir et al., 2023), it is likely that future regulatory policies will demand time-based and location-based reporting of operational emissions. Based on current initiatives from industry (Energy, 2023; Cramer and Goodman, 2023) and legislation (Becker, 2021), we assume that operational carbon emissions of electricity usage will be reported hourly—or even sub-hourly (Haugen et al., 2024; Bashir et al., 2023).
We also follow the predominant assumption in related works that future regulations will use average carbon intensity (ACI), reported in carbon-equivalent GHG per unit of energy (gCO2/kWh), as a fundamental metric to determine the operational carbon emissions at specific times and locations (Gsteiger et al., 2024; Wiesner et al., 2021; Hanafy et al., 2024b; Sukprasert et al., 2024b; Radovanovic et al., 2022). ACI is defined as a weighted average of all produced energy in a region, weighted by the carbon intensity of the energy source. Some organizations, like Microsoft (Buchanan et al., 2023) advocate the use of marginal carbon intensity (MCI), which refers to the additional carbon emissions generated from meeting incremental demand at a given time. Although the two metrics can result in different scheduling decisions (Sukprasert et al., 2024a), both can be used in our framework. In our analysis, however, we assume ACI, as MCI is very ambiguous to compute in practice (Wiesner et al., 2021) and therefore a lot less likely to make it into official policies.
3. Carbon-Aware QoR Adaptation
Offering different service quality tiers is common practice in the operation of cloud services. For example, content delivery networks provide faster and more reliable service to higher-tier users, while AI platforms like OpenAI and Gemini reserve access to larger and more advanced models for paying customers. Furthermore, service providers already degrade service quality during periods of high load to prevent system overload (Park et al., 2024; Meza et al., 2023). Such measures often target best-effort users: For instance, free-tier users of LLM inference services can experience highly increased latencies at certain times, to not endanger the QoS of subscribed users.
Recently, there have been proposals to design applications that can intentionally sacrifice QoS, not only to manage economical interests or load peaks, but as a strategy to improve sustainability (Vitali, 2022; Vitali et al., 2023). Especially in AI inference, we have numerous options for fine-grained quality tiering that have different impacts on service quality and energy efficiency. For instance, service quality tiers for LLMs can vary based on factors such as the choice of model (e.g., number of parameters), quantization (ranging from 4-bit integers to 32-bit floating-point), and the number of tokens generated (i.e. providing shorter replies (Li et al., 2024b)), and can have a significant impact on the carbon footprint of inference.
In this work, we examine the impact of dynamically adjusting service QoR over time to enhance carbon efficiency within a single region, without relying on geo-distributed load balancing.
3.1. Problem Setting
We define an optimization time window $\left[0,T\right]$. Given an interval $\Delta$, we divide the window into $T/\Delta$ time intervals
(1) $\mathcal{T}_{i}=[i\Delta,(i+1)\Delta],\quad i=0,1,\dots,I-1,$
where $I=T/\Delta$ is the total number of intervals, i.e. $i=0$ corresponds to the first interval and $i=I-1$ to the last. For the remainder of this paper, we consider $T=1~{}\text{year}$ and $\Delta=1$ hour.
We define the set of user groups as $\mathcal{U}$, machine types $\mathcal{M}$ and service quality tiers $\mathcal{Q}$. Machine types can represent physical hardware as well as virtual machines (VMs) like cloud instances.
For each time interval, there is an associated carbon intensity $C^{i}$ and a number of requests $R^{i}$ per user group. Each machine can handle a certain number of requests by serving a certain quality tier (denoted as $K$), which corresponds to its current power usage $P^{i}$. The deployment $D^{i}$ must be sufficient to meet the number of incoming requests that are allocated to different service quality tiers ($A^{i}$) [^1] Note, that deployments can be constrained, i.e., $d^{i}_{m,q}\in[0,\dotsc,\bar{d}]$, where $\bar{d}$ represents the maximum number of available machines. Furthermore, when elements in $D^{i}$ are sufficiently large (i.e. scenarios with many active machines), an optimal result can be approximated by relaxing $D^{i}$ to continuous variables. This turns the optimization problem introduced later into a linear program that can be solved in polynomial time.. The key challenge lies is determining the optimal values for both $D^{i}$ and $A^{i}$ for each interval. Table 1 shows an overview of all input and decision variables.
Table 1. Description of symbols. Small letters denote the elements of the corresponding, capitalized matrix.
Input variables
$\mathcal{U}$ Set of user groups $u$
$\mathcal{M}$ Set of machine types $m$
$\mathcal{Q}$ Set of service quality tiers $q$
$\gamma$ Validity period length
$R^{i}\in\mathbb{R}^{\left|\mathcal{U}\right|}$ Number of requests $\left(r^{i}_{u}\right)$ for user group $u$ during  $\mathcal{T}_{i}$.
$C^{i}\in\mathbb{R}$ Carbon intensity during $\mathcal{T}_{i}$.
$P^{i}\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Power usage $\bigl{(}p^{i}_{m,q}\bigl{)}$ of machine $m$ serving quality $q$ during $\mathcal{T}_{i}$.
$K\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Requests per timestep $\bigl{(}k_{m,q}\bigl{)}$ that machine $m$ can serve at quality $q$.
$C_{m}^{\text{emb}}$ Attributed embodied emissions for running machine $m$ for time $\Delta$.
Decision variables
$D^{i}\in\mathbb{N}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Deployment $\bigl{(}d^{i}_{m,q}\bigl{)}$, i.e., the number of machines $m$ running quality $q$ during $\mathcal{T}_{i}$.
$A^{i}\in\mathbb{R}^{\left|\mathcal{U}\right|\times\left|\mathcal{Q}\right|}$ Number of requests $\bigl{(}a^{i}_{u,q}\bigl{)}$ served by quality $q$ for user group $u$ during $\mathcal{T}_{i}$.
3.2. Quality of Responses
We define $\text{QoR}\in[0,1]$ as a metric that describes the proportion of requests served by different service quality tiers. In our subsequent analysis, we concentrate on a single user group, such as best-effort users, whose requests can be served by two distinct service tiers that differ in both response quality and energy demand. In this scenario, $\text{QoR}=0$ indicates that all requests are served by the lower-quality Tier 1, while $\text{QoR}=1$ indicates that all requests are served by the higher-quality but more energy-intensive Tier 2. When $\text{QoR}=0.5$, half of the requests are served by Tier 1 and half by Tier 2.
In the following, we introduce several metrics to establish a more general definition of QoR, that holds for multiple user groups and service quality tiers.
Service level objectives.  Adopting common terminology for quantifying QoS metrics (Keller and Ludwig, 2003), users of this framework can define a lower-bound service level objective $\underline{\text{SLO}}\in[0,1]^{\left|U\right|\times\left|Q\right|}$, where $\underline{\text{\text{slo}}}_{u,q}$ indicates the minimal proportion of requests that should be served by quality tier $q$ for user group $u$. Similarly, users can define an upper-bound $\overline{\text{SLO}}\in[0,1]^{\left|U\right|\times\left|Q\right|}$ to service quality. For example, LLM service providers often allow non-paying users a limited number of requests to access their highest-quality models, but do not intend to serve all requests at this quality tier.
[…]
$\underline{\text{SLO}}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\quad\text{and}\quad\overline{\text{SLO}}=\begin{bmatrix}0&1\\ 0.6&0.4\end{bmatrix}$
describes that premium users’ requests must always be served by Tier 2, while 60–100 % of requests by best-effort users can be served by Tier 1 and 0–40 % by Tier 2.
All requests per user group must be served by a quality tier:
(2) $\sum_{q\in\mathcal{Q}}\underline{\text{slo}}_{u,q}=1\quad\text{and}\quad\sum_{q\in\mathcal{Q}}\overline{\text{slo}}_{u,q}=1\quad\forall u\in\mathcal{U}.$
Service level indicators.  The service level indicator $\text{SLI}(\alpha,\omega)$ describes the observed distribution of requests across different quality tiers within the interval $\bigcup_{i=\alpha,\dotsc,\omega}\mathcal{T}_{i}$:
(3) $\text{sli}_{u,q}(\alpha,\omega)=\frac{\sum_{i=\alpha}^{\omega}a^{i}_{u,q}}{\sum_{i=\alpha}^{\omega}r^{i}_{u}},$
As the service level is constrained by the SLOs, each
(4) $\text{sli}_{u,q}(\alpha,\omega)\in\left[\min(\underline{\text{slo}}_{u,q},\overline{\text{slo}}_{u,q}),\max(\underline{\text{slo}}_{u,q},\overline{\text{slo}}_{u,q})\right].$
General definition of QoR.  We define the service’s quality of responses $\text{QoR}(\alpha,\omega)$ within an interval $\bigcup_{i=\alpha,\dotsc,\omega}\mathcal{T}_{i}$ through the relation between $\text{SLI}(\alpha,\omega)$, $\underline{\text{SLO}}$ and $\overline{\text{SLO}}$. In particular, for the set of parameters that permit optimization, defined as $\mathcal{V}=\{(u,q)\in\mathcal{U}\times\mathcal{Q}:\overline{\text{slo}}_{u,q}\neq\underline{\text{slo}}_{u,q}\}$, we define QoR as:
(5) $\text{QoR}(\alpha,\omega)=1-\min_{(u,q)\in\mathcal{V}}\left(\frac{\big{|}\overline{\text{slo}}_{u,q}-\text{sli}_{u,q}(\alpha,\omega)\big{|}}{\big{|}\overline{\text{slo}}_{u,q}-\underline{\text{slo}}_{u,q}\big{|}}\right).$
Following the above example, an $\text{SLI}(\alpha,\omega)=\begin{bmatrix}0&1\\ 0.7&0.3\end{bmatrix}$ would result in a $\text{QoR}(\alpha,\omega)=0.75$.
Validity periods.  As is typical in the assessment of QoS metrics (Keller and Ludwig, 2003), QoR is defined over validity periods $[\alpha,\alpha+1,...,\omega]$. These periods can either be disjoint (e.g., quarters, months, weeks) or—more commonly—overlapping. For the remainder of this paper, we assume that QoR is assessed over rolling windows of length $\gamma$ (integer multiple of adjustment interval $\Delta$). Staying with the above example, when the validity period is set to $\gamma=24~{}\text{h}$ and we want to guarantee a specific $\text{QoR}_{\text{target}}=0.5$, we have to ensure that within each 24 hour window at least 50 % of all requests are served by Tier 2. Formally, $0.5\leq\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}$.
Figure 3 illustrates how the validity period length influences the hourly QoR over time. Longer validity periods provide greater flexibility in balancing service quality in response to fluctuations in carbon intensity, but can lead to prolonged periods of high or low hourly QoR.
Figure 3. Increasing validity periods allow for more flexibility in adjusting the proportion of requests served by Tier 1 and Tier 2 based on carbon intensity.
3.3. Power Model
To determine the operational carbon emissions from powering servers, we define the power usage $P^{i}=\bigl{(}p^{i}_{m,q}\bigl{)}$ of a machine $m$ running service quality $q$ during interval $\mathcal{T}_{i}$. We describe two power models from the perspective of a data center operator and a cloud tenant that report their Scope 2 and 3 emissions, respectively.
[…]
(6) $p^{i}_{m,q}=\text{PUE}\cdot\left(p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot\left(\text{util}^{i}_{q}\right)^{n}\right),$
where
(7) $\text{util}^{i}_{q}=\frac{\sum_{u\in\mathcal{U}}a^{i}_{u,q}}{\sum_{m\in\mathcal{M}}\left(d^{i}_{m,q}\cdot k_{m,q}\right)}.$
Cloud tenant perspective.  Service providers that rent their infrastructure from cloud providers, indirectly report the cloud data center’s power usage as part of their Scope 3 emissions. Although cloud providers recently started reporting emissions caused by renting VMs, their methods are inconsistent, lack transparency, and are subject to change. Until providers (are required to) adopt a standardized, publicly available methodology, we are left with estimating this factor. We assume that cloud providers attribute a constant, load-independent factor $P_{m}^{\text{attr}}$ to each active machine $m$:
(8) $p^{i}_{m,q}=p_{m}^{\text{attr}}$
3.4. Carbon Emission Model
We define a service’s estimated carbon emissions $E^{i}$ during time period $\mathcal{T}_{i}$ with respect to the region’s current operational carbon intensity $C^{i}$ (in gCO2/kWh) and the machine’s embodied emissions $C_{m}^{\text{ emb}}$ (in gCO2):
(9) $E^{i}=\sum_{m\in\mathcal{M}}\sum_{q\in\mathcal{Q}}d^{i}_{m,q}\cdot\left(p^{i}_{m,q}\cdot C^{i}+C_{m}^{\text{emb}}\right)$
Depending on methodology, embodied emissions might not always be reported with respect to infrastructure usage, in which case $C_{m}^{\text{emb}}=0$.
3.5. Optimization Problem
We formalize the problem of optimizing service deployments $D^{i}$ that minimize carbon emissions, while providing an allocation $A^{i}$ of requests to different service quality tiers that satisfies $\text{QoR}_{\text{target}}$:
(10) $\min\quad$ $\sum_{i=\alpha}^{\omega}E^{i}$
(11) s.t. $\sum_{q\in\mathcal{Q}}a^{i}_{u,q}=r^{i}_{u}$ $\forall i,\forall u$
(12) $\sum_{u\in\mathcal{U}}a^{i}_{u,q}\leq\sum_{m\in\mathcal{M}}d^{i}_{m,q}\cdot k_{m,q}$ $\forall i,\forall q$
[…]
(14) $\text{QoR}_{\text{target}}\leq\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}$
where
  • All requests served (11) ensures that all requests during each interval $\mathcal{T}_{i}$ are attributed to a quality tier.
  • Sufficient resources (12) ensures that the provided machines have sufficient capacity to serve all requests per quality tier.
  • Respect SLOs (13) ensures that all SLOs are respected.
  • Sufficient QoR (14) ensures that the QoR of every validity period stays above $\text{QoR}_{\text{target}}$.
Note, that this model assumes knowledge of future carbon intensity ($C^{i}$) and requests ($R^{i}$) and can therefore not be solved in practice. However, we use it for assessing upper-bound potential and as the foundation of the online approach proposed in Section 4.
3.6. Complexity
The proposed optimization problem is an NP-hard mixed-integer linear program (MILP), as we minimize a linear objective function subject to linear constraints, which include discrete decision variables in $D^{i}$.
Equivalence of power models.  From the perspective of data center operators, the load-dependant power model from (6) introduces a non-linear relationship by multiplying values within $D^{i}$ and $A^{i}$. This would turn the MILP into an mixed-integer quadratic program, which is considerably harder to solve. However, during optimization, we can relax (6) to the same form as (8), as long as $n\leq 1$, i.e. the power model is concave.
Theorem 3.1.
We assume $p_{m,q}^{\text{max}}>p_{m}^{\text{idle}}>0$ and $n\in[0,1]$. When minimizing (10) subject to (11) – (14) from the perspective of data center operators, the utilization-dependent power model,
$p^{i}_{m,q}=p_{m}^{\text{idle}}+\bigl{(}p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\bigl{)}\cdot\bigl{(}\text{util}^{i}_{q}\bigl{)}^{n},$
can be simplified to
$p^{i}_{m,q}=p_{m,q}^{\text{max}},$
as it always yields an equivalent optimal deployment $D^{i}$.
The proof is provided in Appendix A. Intuitively, only if the power-law relationship is strictly convex ($n>1$), it can be more efficient to deploy multiple machines at lower utilization rather than fully utilizing fewer machines. However, this case cannot be observed in cloud data centers, where a main objective for energy efficiency is VM consolidation, to minimize the number of active machines (Xu et al., 2017; Gill and Buyya, 2018; Barroso and Hölzle, 2007). Hence, for $n\leq 1$, we can use a load-independent power model during optimization, as it yields the same optimal $D^{i}$.
4. Online Approach
For assessing the practicability of carbon-aware QoR adaptation under realistic conditions, we propose an online approach, which iteratively determines the optimal $A^{i}$ and $D^{i}$ for every interval $\mathcal{T}_{i}$. Its performance is limited by two factors:
  1. (1)
    the error of request and carbon intensity forecasts
  2. (2)
    the error introduced by approximating the NP-hard optimization problem
In our experiments, conducted with realistic forecasts and strict time limits for solving the MILP, we achieve emission savings of $82\pm 6$ relative to the upper-bound potential (Section 5.4).
4.1. Multi-Horizon Optimization
Based on the optimization problem introduced in Section 3.5, we propose a forecast-based multi-horizon optimization which determines the optimal deployment and load balancing for every $\mathcal{T}_{i}$.
Forecasting.  There exist various approaches for long-term load forecasting, such as Prophet (Taylor and Letham, 2017), DeepAR (Flunkert et al., 2017), or TBATS (De Livera et al., 2011). As the electricity mix of regions does not change abruptly, long-term forecasts for ACI can be usually estimated reasonably well by fitting daily, weekly, and annual seasonalities. However, especially in regions that are heavily dependent on weather-dependent renewable power production, additionally using short-term forecasts can provide significantly higher accuracy (Maji et al., 2023; Li et al., 2024c). For example, CarbonCast (Maji et al., 2023) reports mean absolute percentage errors (MAPE) of only 3-10 % for 24-hour forecasts.
Algorithm.  The approach is described in Algorithm 1. As the optimization problem is computationally hard, we approximate near-optimal solutions by either imposing time limits or terminating the optimization process once it reaches a solution within an acceptable proximity to optimality. To reduce the negative impact of these approximations on our decisions, we perform the optimization in two distinct steps with different levels of precision:
Algorithm 1 Multi-Horizon Optimization
1:for $\alpha\leftarrow 0$ to $I-1$ do
2:    # Long-term optimization
3:   if $\alpha\equiv 0\pmod{\tau}$ then
4:      $(R^{i},C^{i})_{i=\alpha}^{I-1}\leftarrow$ update forecasts
5:      $(D^{i},A^{i})_{i=\alpha}^{I-1}\leftarrow\min_{(D^{i},A^{i})_{i=\alpha}^{I-1}}\sum_{i=\alpha}^{I-1}E^{i}$
6:    # Short-term optimization
7:   $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma}\leftarrow\min_{(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma}}\sum_{i=\alpha}^{I-1}E^{i}$
8:    # Progress in time
[…]
10:   $R^{\alpha},C^{\alpha},D^{\alpha},A^{\alpha}\leftarrow$ update past with observed reality
  1. (1)
    A long-term optimization, executed every $\tau$ intervals (for example, every 24 hours), solves the MILP for the remainder of the year. Any $D^{i}$ and $A^{i}$ in the past are fixed, i.e. we only optimize $(D^{i},A^{i})_{i=\alpha}^{I-1}$, where $\alpha$ is the current interval. In this step, longer execution times and less optimal solutions are acceptable.
  2. (2)
    A short-term optimization, executed every interval, solves the problem within a fixed horizon [^2] We found that the validity period length $\gamma$ is a good indicator for horizon length, although it can be shorter or longer (as future values of $A^{i}$ are fixed by the long-term optimization). In Line 7 we assume $\gamma$ as the horizon.. As we only optimize over $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma}$, this optimization involves significantly fewer variables and is more likely to find an optimal or near-optimal solution quickly. If no solution is found, we select an $A^{i}$ such that $\text{QoR}(\alpha,\omega)=1$, and determine the minimal $D^{i}$ that satisfies the Sufficient resources (12) constraint.
After optimization, the determined deployment $D^{\alpha}$ is provisioned for the current interval. As short-term load and carbon intensity predictions are usually very precise, and the provisioning of new nodes can take significant time (Microsoft reports 6-8 minutes for creating a new LLM inference instance (Stojkovic et al., 2024)) we assume that there are usually no rapid auto-scaling decisions within an interval. After the interval has passed, we update the actually observed $D^{\alpha}$ and $A^{\alpha}$, alongside the observed number of requests and carbon intensity.
4.2. Automatic QoR Adaptation
The optimization problem introduced in Section 3.5 minimizes emissions under a given $\text{QoR}_{\text{target}}$. However, especially for best-effort users, another relevant question is: What are the expected annual emissions from this user group, and can we keep them within a predefined budget?
In Section 5.4, we show how running multiple long-term optimizations can provide service providers with valuable insights into expected annual emissions, allowing for adjustments to $\text{QoR}_{\text{target}}$ throughout the year. Building on this, we propose a variation of the online approach introduced in Section 4, which finds an optimal $\text{QoR}_{\text{target}}$ for best-effort users to stay within a carbon budget $B$:
(15) $\max\quad$ $\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}$
[…]
(16) $\sum_{i=\alpha}^{\omega}E^{i}\leq B$
In the multi-horizon optimization (Algorithm 1), we replace the long-term optimization in Line 5 with this objective:
(17) $(D^{i},A^{i})_{i=\alpha}^{I-1}\leftarrow\max_{(D^{i},A^{i})_{i=\alpha}^{I-1}}\left(\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=0}^{I-1}\right)$
Note, that we solve the optimization over $i=0,\dots,I-1$ (not just $i\geq\alpha$) as we must aggregate emissions over all intervals in (16). This has no performance implication, as we still only optimize the variables $(D^{i},A^{i})_{i=\alpha}^{I-1}$. Past values in $D^{i}$ and $A^{i}$ are fixed.
5. Experiments
We assess the potential of carbon-aware QoR adaptation by simulating the execution of an LLM inference service over the year 2023 using 8 real and synthetic request traces in 10 different regions.
We use Gurobi (Gurobi Optimization, LLC, 2024) to solve optimization problems. All simulations were conducted under the same conditions on an HPC cluster, with 16 cores allocated to each job.
Table 2. Upper-bound potential and carbon savings under realistic conditions in % for all request traces and regions at $\gamma=1$ week and $\text{QoR}_{\text{target}}=0.5$. Note, that other $\text{QoR}_{\text{target}}$ can lead to even better results (Figure X).
Region Artificial Real Synthetic Mean
Static Random Wiki (en) Wiki (de) Taxi Cell B Cell D Cell F
NL 8.2 / 7.4 8.5 / 6.3 7.6 / 6.5 10.2 / 9.3 11.3 / 8.7 7.8 / 6.7 7.6 / 6.6 8.0 / 7.0 8.7±1.3 / 7.3±1.0
CISO 7.0 / 6.8 7.5 / 6.0 7.5 / 7.0 10.0 / 9.4 11.3 / 9.2 6.6 / 6.1 6.3 / 5.9 6.8 / 6.4 7.9±1.7 / 7.1±1.3
ES 7.1 / 6.3 7.3 / 5.2 6.5 / 5.6 9.4 / 8.4 10.4 / 8.0 6.6 / 5.6 6.6 / 5.7 6.9 / 5.9 7.6±1.4 / 6.3±1.1
AU-QLD 7.9 / 7.8 8.4 / 7.0 7.1 / 6.5 6.1 / 5.8 8.8 / 6.9 7.6 / 7.0 6.9 / 6.4 8.0 / 7.5 7.6±0.8 / 6.9±0.6
DE 6.3 / 5.4 6.7 / 4.5 5.7 / 4.6 8.7 / 7.6 9.7 / 6.9 5.9 / 4.8 5.8 / 4.8 6.1 / 5.1 6.9±1.4 / 5.5±1.1
PL 4.1 / 3.8 4.5 / 2.9 3.5 / 2.9 6.7 / 6.2 7.5 / 5.7 3.8 / 3.3 3.7 / 3.2 4.1 / 3.5 4.7±1.4 / 3.9±1.2
ERCOT 4.2 / 3.1 4.6 / 2.6 3.6 / 2.4 6.3 / 5.1 7.4 / 5.8 3.9 / 2.8 3.7 / 2.5 4.1 / 3.0 4.7±1.3 / 3.4±1.2
SE 2.7 / 2.4 3.0 / 1.6 2.0 / 1.5 4.9 / 4.5 6.1 / 5.4 2.3 / 1.9 2.2 / 1.7 2.6 / 2.2 3.2±1.4 / 2.6±1.4
NYISO 2.3 / 1.8 2.7 / 1.1 1.7 / 1.1 4.6 / 4.3 5.6 / 4.6 1.9 / 1.5 1.8 / 1.3 2.2 / 1.7 2.8±1.4 / 2.2±1.3
PJM 1.8 / 1.7 2.3 / 0.8 1.3 / 1.0 4.1 / 4.0 5.5 / 4.3 1.6 / 1.4 1.4 / 1.1 1.8 / 1.6 2.5±1.4 / 2.0±1.3
5.1. Experimental Setup
As ACI is usually reported hourly, we define $\Delta=1$ hour.
[…]
Request traces.  For our analysis, we require 4 years of hourly request traces: 2020-2022 for fitting forecasting models and 2023 for the analysis itself. As such data are not publicly available for LLM inference servies, we show the effectiveness of our method across a diverse set of other real and synthetic request traces:
  • Artificial: A Static trace that assumes a constant stream of hourly requests, and a Random trace, where we sample hourly requests from a normal distribution.
  • Real: Two real-life traces representing the English and German Wikipedia Pageview statistics (Wikimedia Foundation, [n. d.]), Wiki (en) and Wiki (de), and a trace based on the hourly taxi trips in New York City (Taxi and Commission, 2025), called Taxi.
  • Synthetic: Three generated traces, called Cell B, D, and F, which are based on instance event traces of Borg cluster cells at Google (Google, 2019).
All request trace datasets and our forecast methodology are explained in detail in Appendix B.
Carbon intensity traces.  We evaluate the potential for ten globally distributed regions:
  • In Europe, we consider Germany (DE), Spain (ES), Netherlands (NL), Poland (PL), and Sweden (SE)
  • In the US, we consider California (CISO), Texas (ERCOT), New York (NYISO), and the Pennsylvania-Jersey-Maryland Interconnection (PJM)
  • In Australia, we consider Queensland (AU-QLD)
All carbon intensity data was provided by ElectricityMaps (Maps, 2024). We normalized all time zones to align the time of day with the request traces. The temporal variability of regions is depicted in Figure 1. Our forecast methodology is explained in Appendix C.
5.2. Absolute Emissions
We first analyze the absolute annual emissions without carbon-aware QoR adaption, i.e. for $\gamma=1$ hour. The results reflect emission savings driven purely by increased energy demand from serving more requests on Llama 3.1 70B ($\text{QoR}_{\text{target}}\to 1$). Figure 4 shows the annual emissions for Wiki (en)—the trend is equivalent for all datasets. We observe very large differences across different regions: For example, Sweden’s energy mix relies almost entirely on renewable energy and nuclear power, making it one of the least carbon-intensive regions globally. Serving all requests with the 70B model results in just 245 tCO2, almost 27 times less than in Poland (4681 tCO2), where coal is the dominant energy source. As expected, the resulting emissions scale linearly with $\text{QoR}_{\text{target}}$.
Figure 4. Annual operational carbon emissions for Wiki (en) at different $\text{QoR}_{\text{target}}$ without carbon-aware QoR adaption.
Figure 5. Additional carbon savings resulting from increased validity periods, relative (top) and absolute (bottom) for $\text{QoR}_{\text{target}}=0.5$. Bigger validity periods yield larger gains.
5.3. Upper-Bound Potential
[…]
Carbon savings beyond energy efficiency.  In blue (left), Table 2 presents the relative additional carbon savings for $\text{QoR}_{\text{target}}=0.5$ at $\gamma=1$ week compared to $\gamma=1$ hour, where emissions directly depend on energy demand. We observe that through carbon-aware QoR adaptation, emissions drop by around 8 % in many regions. While request trace patterns exhibit little difference in potential, traces with a low volume of requests (Wiki (de) and Taxi) show increased potential, due to the higher impact of discrete scheduling decisions. For example, the volume of requests in Wiki (de) results in 10–25 active machines at a time, while Wiki (en) requires 80–200 active machines. In experiments with even fewer machines we managed to gain relative savings of more than 40 %—however, we do not analyze such cases further, as we focus on high-impact scenarios with many machines.
Impact of validity periods.  Figure 5 reports the relative and absolute savings for $\text{QoR}_{\text{target}}=0.5$ and different validity periods. Increasing $\gamma$ to 8 hours yields only limited additional savings—less than 3 % across all settings. This is because average carbon intensity shows limited short-term variation, leaving little room for optimization. However, starting at $\gamma=24$ hours, we see a more notable potential of 5-8 % in some regions. Regional characteristics significantly influence the expected gains: For instance, while California’s carbon intensity primarily fluctuates on a daily basis due to its heavy reliance on solar energy, Germany exhibits more complex patterns with daily, weekly, and seasonal variations (Wiesner et al., 2021). As a result, extending $\gamma$ beyond 24 hours has minimal effect in California, while in Germany, larger $\gamma$ values lead to considerable improvements. In contrast, regions like Sweden, New York, and PJM exhibit very little temporal variation in carbon intensity.
Impact of premium users.  Figure 6 illustrates the loss of potential with a growing fraction of premium users at $\gamma=1$ week and $\text{QoR}_{\text{target}}=0.5$. We observe a slightly sub-linear decrease with a certain offset which depends on the overall request volume and, hence, the impact of discrete scheduling decisions. For reference, ChatGPT reached over 200 million active users in 2024 (Reuters, 2024), with fewer than 10 million paying customers (Tarantola, 2024). The resulting 5 % premium users decrease the potential savings by 5.6 % on average.
Figure 6. Impact of premium users on the potential savings.
Impact of $\text{QoR}_{\text{target}}$.  Figure 7 shows the potential for relative savings across different $\text{QoR}_{\text{target}}$. At $\text{QoR}_{\text{target}}=0$, we serve all requests using LLaMA 3.1 8B, leaving no flexibility for additional savings. Likewise, when $\text{QoR}_{\text{target}}$ is 1, all requests must be served by LLaMA 3.1 70B. A $\text{QoR}_{\text{target}}$ around 0.5 offers the greatest flexibility for carbon-aware QoR adaptation, resulting in the highest potential for carbon savings.
Figure 7. Additional relative savings for different $\text{QoR}_{\text{target}}$.
Figure 8. CDF for Wiki (en) in Germany at $\text{QoR}_{\text{target}}=0.5$: What proportion of time intervals of length $\beta=\{1,3,7\}$ days have a QoR below the threshold?
Periods of low service quality.  While increasing $\gamma$ enhances potential for carbon savings, it also increases the probability of prolonged periods with low QoR, illustrated in Figure 8. For example, when optimizing for $\text{QoR}_{\text{target}}=0.5$ at $\gamma=1$ week, no 1-week interval has a QoR below 0.5 (right). However, 10 % of all daily intervals exhibit a $\text{QoR}_{\text{target}}<0.48$ (left). We observe that for $\gamma=3$ months, users can experience multiple consecutive days of low QoR. Consequently, very large validity periods of more than a month are not practical in real-world scenarios.
5.4. Evaluation under Realistic Conditions
We evaluated the practicability of carbon-aware QoR adaptation using the proposed multi-horizon optimization under real forecasts (see Appendix B as well as C) and approximated MILP solutions: Long-term optimizations are stopped at a $\text{MILP}_{\text{gap}}<0.1\,\%$ or after 30s; short-term optimizations at a $\text{MILP}_{\text{gap}}<0.1\,\%$ or after 10s.
[…]
Figure 9. Additional carbon savings under realistic conditions for Wiki (de) across all regions and $\text{QoR}_{\text{target}}=0.5$.
Outlook of annual emissions.  Long-term optimizations are computationally cheap when accepting near-optimal solutions—within just 30 seconds, the mean $\text{MILP}_{\text{gap}}$ reached 0.53 % accross all experiments. This allows us to generate a broad spectrum of solutions for varying $\text{QoR}_{\text{target}}$ and request forecasts, to represent, for instance, worst-case or best-case scenarios. Figure 10 demonstrates emission outlooks for Wiki (en) traces in California. For simulating best-case and worst-case request forecasts, we use the lower and upper bound of the 80 % prediction intervals provided by Prophet, respectively. Each week, we solve 33 long-term optimizations, covering the three forecasts and $\text{QoR}_{\text{target}}=\{0,0.1,...,1\}$.
(a) Annual emission outlook at three discrete points in time.
(b) Annual emission outlook; weekly for $\text{QoR}_{\text{target}}=0.5$.
Figure 10. Projected emissions (blue) based on long-term forecasts. The interval shows the range between best- and worst-case forecasts. Red indicates emissions already incurred.
Figure 10(a) shows the resulting outlooks a the start of the year, after 10 weeks, and after 40 weeks. The outlooks provide a comprehensive picture of the future, based on the currently available information. From the initial outlook (left), service providers may conclude that setting $\text{QoR}_{\text{target}}=0.5$ is a reasonable target to stay within an annual carbon budget of 1000 tCO2. Consequently, we optimize for $\text{QoR}_{\text{target}}=0.5$ for the remainder of the year.
Figure 10(b) shows that the initial estimate was overstated, although still in the range of the lower bound It depicts the weekly estimates for annual emissions when aiming for $\text{QoR}_{\text{target}}=0.5$. The upper and lower bound, defined by the worst-case and best-case forecasts, narrow over time as incurred emissions accumulate.
5.5. Automatic QoR Adaptation
Figure 11. Daily QoR for California (CISO), Texas (ERCOT), and Poland (PL). Our approach provides the most consistent QoR throughout the year.
Lastly, we illustrate the automatic adaptation of $\text{QoR}_{\text{target}}$ introduced in Section 4.2 on Wiki (en) in three exemplary regions. For each region, we define an annual budget based on the resulting emissions of the corresponding upper-bound experiment (Section 5.3) for $\gamma=24$ hours and $\text{QoR}_{\text{target}}=0.5$. Consequently, under perfect forecasts, the optimal approach would be to choose $\text{QoR}_{\text{target}}=0.5$ for every time step. We compare our approach to two greedy baselines that have access to the same forecasts as our approach:
  1. (1)
    Greedy (constant) divides the annual budget into equally sized hourly budgets for which it solves (15). Any remaining or overspent budget is evenly distributed across the remaining time intervals.
  2. (2)
    Greedy (weighted) also solves (15), but weights the hourly budgets based on forecasted load and carbon intensity. Every 24 hours, as request and carbon intensity forecasts are updated, weightings get adjusted.
Figure 11 shows the daily QoR of the greedy baselines and our automatic QoR adaptation. Since Greedy (constant) does not account for fluctuations in requests or carbon intensity, the provided QoR varies significantly throughout the year. Greedy (weighted) delivers more consistent QoR in all experiments but tends to fall back to either $\text{QoR}=1$ (NL) or $\text{QoR}=0$ (CISO and NYISO) toward the end of the year due to forecast errors. In the CISO and NYISO regions, it even exceeds the budget by 0.8 and 1.2 %, respectively. Our approach provides the most consistent service across all experiments as it periodically adjusts $\text{QoR}_{\text{target}}$ based on long-term optimizations. The annual budget is utilized to 98.2–99.8 %.
6. Discussion
This section highlights the impact of our results on energy-intensive services and discusses the limitations of this study.
Impact.  We emphasize the significant potential for carbon efficiency improvements in cloud services, based on our analysis in Section 5.2. ChatGPT, one of the most energy-intensive cloud services today, reached over 200 million active users in 2024 (Reuters, 2024). Assuming just 5 requests per user per day, this amounts to nearly 42 million requests per hour on average, which is an order of magnitude more than what we consider in our larger scenarios. Assuming performance similar to Llama 3.1 70B, we estimate that the annual operational carbon emissions of large-scale LLM service providers currently exceed multiple 10,000 tCO2—a figure set to rise with multi-modal AI models.
Furthermore, this study assumes ACI as the most likely metric to be implemented in official policy. However, and increasing part of the community is advocating to base load shifting decisions based on marginal signals (WattTime, 2022; Buchanan et al., 2023; Dodge et al., 2022), which—if implemented properly—better reflects the underlying complexities of electric grids. Yet, to this day, existing methodologies for MCI are based on predicting the marginal generator, which is ambiguous and error prone. Moreover, current reporting lacks spatial granularity to capture effects like energy curtailments from local grid congestion. We predict that with improved methodologies and more granular MCI reporting, carbon-aware QoR adaptation could have much greater potential, as MCI shows higher temporal variability than ACI, enabling the use of shorter validity periods. Our approach is directly applicable to MCI signals or any other scalar metric without requiring adaptations.
Limitations and future work.  This paper presents the first analysis on the potential of carbon-aware QoR adaptation in services, focusing on an LLM inference use case. While our approach is already applicable to more complex scenarios with multiple service quality tiers and heterogeneous infrastructure, such cases have not yet been investigated experimentally.
Second, our current modeling only considers the QoR for user groups, rather than individual users. This assumes a uniform distribution of users within each group and may lead to fairness concerns if the distribution is uneven. For example, ACI tends to peak in the evenings in most regions, meaning that users who usually access the service during these hours can experience a lower QoR than other users in their group. We consider fairness aspects with more granular user modeling as an area for future work.
Third, future work should explore the impact of adaptive QoR adjustments on user quality of experience and the resulting implications on user behavior. For instance, in the context of generative AI, reduced QoR might incentivize users to resubmit requests, potentially offsetting carbon savings. Unfortunately, to the best of our knowledge, there are no established behavioral models for generative AI usage. Developing such models would be valuable to better quantify the real-life impacts of sustainability approaches, including rebound effects.
7. Related Work
We survey related works on carbon-aware computing as well as carbon efficiency in energy-intensive cloud services.
Carbon-aware cloud services.  The large majority of approaches in carbon-aware computing focuses on shifting flexible batch processing towards times (Wiesner et al., 2021; Radovanovic et al., 2022; Fridgen et al., 2021; Lechowicz et al., 2023) and locations (Zheng et al., 2020; Lindberg et al., 2022; Wiesner et al., 2024) with clean energy. Hence, for batch jobs, the limitations and trade-offs of carbon-aware spatio-temporal workload shifting have been investigated thoroughly (Hanafy et al., 2024b; Sukprasert et al., 2024b; Hanafy et al., 2024a; Lechowicz et al., 2024). However, carbon efficiency in continuously running services that usually have little to no delay-tolerance has received only little attention so far.
Existing approaches for carbon-aware services primarily focus on load balancing across geo-distributed data centers and the respective capacity right-sizing under SLA constraints (Zhou et al., 2013, 2016). For example, Casper (Souza et al., 2023) is a carbon-aware scheduling and provisioning system for web services that minimizes the carbon footprint under latency constraints. CDN-Shifter (Murillo et al., 2024) explores carbon-aware spatial load shifting for content delivery networks. Caribou (Gsteiger et al., 2024) offloads serverless workflows across geo-distributed regions based on average carbon intensity. However, there are no existing approaches that address services which must operate in a single region due to data, security, or infrastructure constraints.
Sustainability in generative AI inference.  Although the carbon footprint of AI training has been an active topic of research for some time (Patterson et al., 2021; Dodge et al., 2022; Strubell et al., 2020; Dhar, 2020) only recently there has been an increasing focus on AI inference (Chien et al., 2023; Eilam et al., 2023; Samsi et al., 2023). Especially, the rapid adoption of LLMs has resulted in multiple studies that try to quantify their growing carbon footprint (Nguyen et al., 2024; Faiz et al., 2024; Luccioni et al., 2024). Many recent approaches target sustainability aspects of inference services throuh power-aware scheduling (Stojkovic et al., 2024, 2025; Qiu et al., 2025; Patel et al., 2024) or by considering their carbon and water footprint in geographic load balancing (Li et al., 2024d).
Service quality adaptation.  While service degradation is a common strategy to prevent system overload (Meza et al., 2023; Park et al., 2024), recent research has proposed designing applications with different service quality tiers as a means of improving sustainability (Vitali, 2022). In the context of carbon-aware LLMs, we identified only one approach, called Sprout (Li et al., 2024b), which optimizes the autoregressive generation process. Sprout returns shorter responses during periods of high carbon intensity, as the energy demand directly correlates with the number of generated tokens. Thereby, it effectively adapts QoR over time and can be further enhanced by leveraging the optimized infrastructure deployments and load balancing proposed in this paper.
8. Conclusion
As data centers and cloud computing infrastructure continue to expand, their role in emissions reduction will be crucial for meeting global climate targets. In this paper, we explored the potential of adjusting a cloud service’s QoR to improve both energy and carbon efficiency. Our findings demonstrate that, in many regions, carbon-aware QoR adaptations can reduce emissions by up to 10 %, beyond the savings from reduced energy demand. Moreover, we propose an approach to automatically adapt the QoR of best-effort users to stay within a predefined annual carbon budget, which is an important contribution for service providers facing increasing pressure from carbon reporting and pricing.
Acknowledgements.
This research was supported by the German Ministry for Education and Research (BMBF) as BIFOLD—Berlin Institute for the Foundations of Learning and Data (ref. 01IS18025A and 01IS18037A).
Appendix A Proof of Theorem 3.1
Theorem 3.1 We assume $p_{m,q}^{\text{max}}>p_{m}^{\text{idle}}>0$ and $n\in[0,1]$. When minimizing (10) subject to (11) – (14) from the perspective of data center operators, the utilization-dependent power model,
$p^{i}_{m,q}=p_{m}^{\text{idle}}+\bigl{(}p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\bigl{)}\cdot\bigl{(}\text{util}^{i}_{q}\bigl{)}^{n},$
can be simplified to
$p^{i}_{m,q}=p_{m,q}^{\text{max}},$
as it always yields an equivalent optimal deployment $D^{i}$.
Lemma A.1.
Increasing the utilization of machines does not result in additional penalties, so there is no incentive to deploy more machines than strictly necessary. Formally, if $p_{m,q}^{\text{max}}>p_{m}^{\text{idle}}>0$ and $n\in[0,1]$, an utilization-dependent power model $f$ is strictly subadditive
(18) $f(x+y)<f(x)+f(y),$
where $x$ and $y$ represent utilization levels $\text{util}^{i}_{q}\in(0,1]$.
Proof of Lemma A.1.
We show that $f$ is strictly subadditive for the two cases $n=1$ and $n<1$.
Case 1: Linear power model ($n=1$):
(19) $f(x+y)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot\left(x+y\right)$
$f(x+y)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)x$
(20) $\quad+p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)y-p_{m}^{\text{idle}}$
(21) $f(x+y)$ $=f(x)+f(y)-p_{m}^{\text{idle}}$
(22) $f(x+y)$ $<f(x)+f(y)$
Case 2: Sub-linear power model ($n<1$): For a strictly concave function $f$ with $f(0)\geq 0$, $f$ is strictly subadditive on $(0,\infty)$.
(23) $f(x)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot x^{n}$
(24) $f^{\prime}(x)$ $=n\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot x^{n-1}$
(25) $f^{\prime\prime}(x)$ $=\underbrace{n(n-1)\vphantom{\Bigl{(}}}_{<1}\ \underbrace{\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)}_{>1}\cdot\underbrace{x^{n-2}\vphantom{\Bigl{(}}}_{>1}<0$
(26) $f^{\prime\prime}(x)<0$ $\implies f\text{ is strictly concave}$
Additionally, $f(0)=p_{m,q}^{\text{max}}>0$. ∎
Proof of Theorem 3.1.
For each interval $i$, we define the minimal deployment $\check{D}^{i}$ as any feasible solution that uses the smallest possible number of machines:
[…]
  1. (1)
    Feasibility with fewer machines is impossible. If a solution $\widehat{D}^{i}$ has fewer machines than $\check{D}^{i}$ in at least one dimension, it would violate Sufficient resources (12).
  2. (2)
    Using extra machines is suboptimal. By Lemma A.1, under the utilization-dependent power model, there is no incentive to deploy extra machines. Likewise, any extra machine increases the objective under the simplified power model.
We conclude that, for both power models, any optimal solution must yield an equivalent deployment $\check{D}^{i}$ for each interval $i$. ∎
Appendix B Request Traces
This section describes all request trace datasets used in the experiments. The traces are visualized in Figure 12; Table 3 presents summary statistics.
Table 3. Request trace and 24-hour forecast statistics.
Dataset statistics ($1\times 10^{6}$) Forecast statistics
Mean ± Std Dev Min Max MAPE
Static 1.00 ± 0.00 1.00 1.00 0.0 ± 0.0
Normal 1.00 ± 0.34 0.00 2.36 38.6 ± 24.6
Wiki (en) 3.38 ± 0.80 1.88 16.41 13.9 ± 8.4
Wiki (de) 0.42 ± 0.24 0.04 1.56 32.1 ± 15.3
Taxi 0.33 ± 0.14 0.04 0.71 26.5 ± 7.1
Cell B 1.94 ± 0.61 0.73 4.10 27.2 ± 13.5
Cell D 2.87 ± 0.80 1.02 7.76 22.1 ± 15.8
Cell F 1.58 ± 0.41 0.87 4.32 18.2 ± 9.3
Figure 12. First month of data and KDE over all data points in 2023 for the eight datasets used in the evaluation.
Artificial Datasets.  We artificially create two traces:
  • Static assumes a constant stream of $1\times 10^{6}$ hourly requests.
  • Random comprises points randomly sampled from a normal distribution with a mean of $\gamma=1\times 10^{6}$ and a standard deviation of $\sigma=0.33\times 10^{6}$.”
During online optimization, we always assume the mean of $1\times 10^{6}$ requests per hour as forecasts.
Real Datasets.  We create three traces based on real data:
  • Wiki (en) is based on the Wikipedia pageview statistics (Wikimedia Foundation, [n. d.]), namely all hourly requests to en.wikipedia.org.
  • Wiki (de) represents all requests to de.wikipedia.org. The trace differs notably from Wiki (en), as the German Wikipedia is primarily accessed from a single timezone, while the English Wikipedia sees traffic from around the globe.
  • Taxi is based on hourly aggregates of taxi trips in New York City (Taxi and Commission, 2025). We sum over all trip records (Yellow Taxi, Green Taxi, For-Hire Vehicle, High Volume For-Hire Vehicle) and multiply the events by factor 10, to bring them into a similar range with the other datasets.
To simulate realistic forecasts, we are using using Prophet (Taylor and Letham, 2017), a state-of-the-art forecasting model that fits daily, weekly, and annual seasonalities. In particular, we fit a model every day at midnight using 3 years of historical data to forecast the rest of the year. The predictions exhibit realistic errors, denoted in Table 3.
Synthetic Datasets.  As all real datasets exhibit very periodic patterns, we additionally created three synthetic datasets to demonstrate that carbon-aware QoR adaptation is feasible on highly unpredictable traces. For this, we aggregated the instance events of the Google cluster traces (version 3) (Google, 2019) individually for the eight available Borg cells. Data was retrieved via the GoogleSQL query:
SELECT
  TIMESTAMP_TRUNC(TIMESTAMP_MICROS(time), HOUR) AS hour,
  COUNT(*) AS num_events
FROM ‘google.com:google-cluster-data.<TABLE>‘
-- Exclude invalid values
WHERE time > 0 AND time < 9223372036854775807
GROUP BY hour
ORDER BY hour;
[…]
We identified the three traces with the lowest 24-hour autocorrelation coefficients, i.e. the traces that exhibit the least daily seasonality: Cell B (0.17), Cell D (0.27), and Cell F (0.22). Since the Google cluster dataset only covers one month of data, we fit a DoppelGANger (Lin et al., 2019) model for each of the traces, which is a state-of-the-art GAN for timeseries generation, and generate 4 years of synthetic data per cell. We apply the same forecasting methodology that is used for real datasets.
Appendix C Carbon Intensity Forecasts
For long-term carbon intensity forecasts, we apply the same forecasting methodology that is used for the request trace datasets. As in (Lechowicz et al., 2024), we generate synthetic short-term forecasts of up to 4 days by adding Gaussian noise to the actually observed carbon intensity data to match the MAPEs reported by the state-of-the-art forecasting model CarbonCast (Maji et al., 2023) for each region, see Table 4. Short-term forecasts are updated daily at midnight.
Table 4. Carbon forecast MAPE over 96 hours as reported in CarbonCast (Maji et al., 2023)
Region Day 1 Day 2 Day 3 Day 4
CISO 8.08 11.19 12.93 13.62
PJM 3.69 4.93 5.87 6.67
ERCOT 9.78 10.93 11.61 12.23
NYISO 6.91 9.06 9.95 10.42
SE 4.29 5.64 6.43 6.74
DE 7.81 10.69 12.80 15.55
PL 3.12 4.14 4.72 5.50
ES 10.12 16.00 19.37 21.12
NL 6.06 7.87 9.08 9.99
AU-QLD 3.93 3.98 4.06 5.87